# D4C: Improving Negative Example Quality to Enhance Machine Abstract Reasoning Ability

### v14

@@ Abstract
This paper is dedicated to addressing the challenge of enhancing the abstract reasoning capabilities of deep learning algorithms, particularly for tasks involving complex human concepts. We introduce Lico-Net, a novel reasoning engine grounded in fuzzy system theory, which encodes the logical structure of Raven’s Progressive Matrices (RPM) problems into fuzzy representations. Lico-Net excels in solving RPM tasks. Furthermore, we propose Lico-Net-Bongard, a tailored version of Lico-Net for the Bongard-Logo problem, which also achieves high reasoning accuracy through fuzzy representations. However, we observe a mismatch between the way deep learning algorithms and humans induce reasoning concepts, primarily attributed to the inadequate quality of negative samples. Improper configuration of negative samples can convey erroneous conceptual information to deep learning algorithms, thereby distorting their learning objectives. To address this issue, we propose two novel approaches: first, treating different sample points within reasoning problems as mutual negative samples to alter the existing negative sample structure in the data; second, designing a negative sample generator based on a step-wise linear attention mechanism to produce high-quality negative samples. Experimental results demonstrate that these methods significantly improve the performance of Lico-Net (-Bongard) and other baseline models on the RPM and Bongard-Logo datasets, as well as in the domain of foundational vision model processing, particularly when addressing the NICO dataset’s distribution shift problem. Our findings emphasize the importance of improving negative sample quality for enhancing the abstract reasoning capabilities of deep learning algorithms and suggest that fuzzy systems represent a promising direction for future research in this field.
@@
Abstract Reasoning, Raven’s Progressive Matrices, Bongard-logo, NICO, Non-Independent and Identically Distributed, Adversarial Learning, Fuzzy System.
@@ I Introduction
Deep learning, as an artificial intelligence technology that mimics the human brain’s operating mechanism, has shone brightly in numerous fields. In generative tasks, it has demonstrated extraordinary talent in image generation [9], text creation, and video production through cutting-edge technologies such as Generative Adversarial Networks (GANs) [7], Variational Autoencoders (VAEs) [8], autoregressive models, and Transformer [4]. Meanwhile, deep neural networks have also achieved remarkable success in various domains, including computer vision [1, 2, 3], natural language processing [4, 5, 6], generative model construction [7, 8, 9], visual question answering systems [10, 11], and visual abstract reasoning [16, 12, 13]. However, deep learning faces challenges in discriminative tasks, which require models to accurately classify or regress based on input data, such as image classification, text categorization, and speech recognition. Despite some achievements in these tasks, deep learning still struggles to deeply understand the human concepts embedded in the data.
This dilemma may stem from deep learning models’ overreliance on superficial statistical patterns in training data, neglecting the deeper structures and logical relationships that underlie human concepts. Therefore, despite the widespread success of deep learning, we still have a long way to go in truly understanding and simulating complex and vague human concepts. This limitation is particularly evident when dealing with data that contains highly complex and ambiguous human concepts.
Given this, our research will focus on discriminative problems involving complex and deep categorical concepts. We will delve into the root causes of deep learning models’ poor performance on such issues and actively explore effective solutions, aiming to drive new breakthroughs in deep learning’s ability to understand and simulate human concepts.
@@ II Problem description
Traditional visual discriminative tasks evaluate the generalization and feature recognition capabilities of deep learning algorithms in terms of pixel arrangements within image data. However, in this section, we introduce several advanced visual discriminative tasks that deviate from the norm. These tasks effectively gauge the ability of deep learning algorithms to learn complex and abstract patterns underlying the visible features. This paper presents a series of fuzzy systems that have excelled in these challenging tasks, highlighting the significance of fuzzy systems as a crucial branch of artificial intelligence.
@@ II-A Raven’s Progressive Matrices
RPM problems, or Raven’s Progressive Matrices, are an authoritative benchmark for assessing the advanced discriminative capabilities of deep learning algorithms due to their requirement for abstract reasoning, pattern recognition, and problem-solving skills. As a high-level, non-verbal task that demands the ability to generalize and transfer knowledge, RPM challenges algorithms to demonstrate cognitive functions akin to human intelligence. Its standardized nature and strong correlation with human intelligence measures make it a relevant and consistent test for evaluating the sophistication of machine learning models, emphasizing its authority in gauging the intelligent level of deep learning algorithms.
@@ II-A1 RAVEN and PGM
The RAVEN database is a collection of Ravens Progressive Matrices problems, each consisting of 16 images. Eight of these images, known as the “statement,” are the main challenge, while the other eight form a pool for selection. The goal is to choose the right images to complete a $3\times 3$ matrix with a progressive geometric pattern, following an abstract ”rule.” The left side of Figure 1 shows an example of RAVEN. Similarly, PGM provides eight images for the stem and answers, but its ”rule” applies to both rows and columns. The right side of Figure 1 gives a typical example of PGM.
@@ II-B Bongard problem
The Bongard problem[12, 13] is a challenging abstract visual reasoning task that assesses pattern recognition by presenting images following one of two contrasting rules. It requires a limited number of training examples and demands the ability to abstract and learn from minimal data, posing a significant challenge for both humans and machine learning algorithms. Despite being a well-recognized test in AI and cognitive science, there is limited literature on automatically solving Bongard problems, which involve not just pattern recognition but also higher cognitive functions like abstract reasoning and concept learning, where current machine learning methods significantly trail human capabilities.
@@ II-B1 Bongrad-logo
The Bongard-Logo[13] problem extends the Bongard problem, serving as a benchmark for human-level concept learning and reasoning. It aims to capture three key aspects of human cognition: context-dependent perception, analogical perception, and perception with limited samples but vast vocabulary. Described using action-oriented LOGO language, Bongard-Logo problems can generate thousands of instances via program-guided techniques. They challenge standard image classification methods’ objectivism and metaphysical realism, necessitating a new, human-like perception. Experimental results show that advanced deep learning methods fall short of human performance, indicating they miss core human cognitive characteristics. Bongard-Logo problems guide research towards a general visual reasoning architecture, aiming to surpass this benchmark.
@@
Bongard-logo problems can be categorized into three parts based on concept domains: 1) Free-form problems (FF), featuring shapes made of random action strokes; 2) Basic shape problems (BA), recognizing pattern of shape combinations; 3) High-level concept problems (HD), assessing abstract concept reasoning, like convexity and symmetry.
@@ III-A RPM solver
Image reasoning models like CoPINet[22], LEN+teacher[21], and DCNet[23] emphasize learning disparities and rules, while NCD[25], SCL[26], and SAVIR-T[27], along with neural symbolism systems (PrAE, NVSA, and ALANS[28, 29, 30]), enhance interpretability and accuracy. RS-CNN and RS-TRAN[31] have particularly excelled at solving RPM problems. Triple-CFN[37] distinguishes itself by implicitly extracting and indexing concept and reasoning unit information, leading to significant reasoning accuracy. CRAB[32], based on Bayesian modeling, has made significant advancements with a customized conservatory for the RAVEN database, despite not tackling core RAVEN challenges, and the scientific community eagerly awaits the broader impact of such innovative methods.
@@ III-B Bongard solver
Addressing Bongard problems, approaches include language feature models, CNNs, and synthetic data. Depweg et al.[12] used Bayesian inference with formal language but faced scalability issues. Kharagorgiev and Yun[35] focused on feature extraction, and Nie et al.[13] experimented with CNNs and meta-learning with limited success. PMoC[36] offers a probabilistic approach to assess group sample likelihoods, proving effective. Triple-CFN[37] stands out for its consistent effectiveness on both Bongard-Logo and RPM problems.
@@ III-C NICO solver
Stable learning [33] addresses the O.o.D. problem by sample reweighting to correct domain distribution shifts and reduce feature correlations, enhancing model performance on unseen test sets. DecAug [34] tackles the two-dimensional O.o.D. problem in NICO data by introducing domain labels, splitting the deep learning model into branches for domain and object learning, with perpendicular parameter learning directions to ensure information decoupling, albeit increasing the workload and categorization difficulty. SimCLR [43] improves representation learning through data augmentation, a learnable nonlinear projector, and the InfoNCE loss function. VICReg [42] introduces three regularization terms on embedded variance for self-monitored image representation learning.
@@ IV Methodology
In this study, we emphasize that deep learning algorithms should focus on learning the fuzzy representations of complex and tightly intertwined human concepts when addressing such issues. By leveraging these fuzzy representations to solve problems, the performance of the algorithms can be significantly enhanced. However, we have observed that, despite the performance boost brought by the introduction of fuzzy representations, there exists a notable discrepancy between the way deep learning algorithms summarize concepts and logical patterns and that of humans. Upon thorough analysis, we found that this mismatch is not attributable to the fuzzy representation technique itself but is a widespread issue in traditional deep learning algorithms. Ultimately, one of the core reasons lies in the inadequate configuration of negative samples in the data. Improper negative sample configuration can convey erroneous conceptual information to deep learning algorithms, thereby distorting their learning objectives.
Therefore, this study posits that, in addition to striving to develop deep learning algorithms capable of effectively learning from low-quality negative samples, improving the quality of negative samples is also a crucial strategy.
@@ V Discriminator of Concepts (DC)
In this section, we introduce Lico-Net, a novel deep neural network-based inference engine designed to efficiently solve RPM (Raven’s Progressive Matrices) problems. By leveraging advanced techniques, Lico-Net fuzzifies the concepts and reasoning logic embedded in RPM problems, achieving excellent performance. Subsequently, we present Loco-Net-Bongard, a deep learning inference engine tailored for solving Bongard-Logo problems, which also attains outstanding reasoning accuracy through explicit concept fuzzification. Furthermore, preliminary experiments on Lico-Net (and its Bongard counterpart) reveal a phenomenon of concept mismatch, highlighting the models’ limitations in learning implicit concepts. Additionally, using the NICO dataset, we further demonstrate how shifts in image background distributions can obscure deep learning models’ ability to learn implicit concepts.
@@ V-A A novel baseline, Lico-Net, for RPM problems
The RPM problem requires participants to select answers from an option pool in order to complete a $3\times 3$ progressive matrix of images. If the completed matrix adheres to a specific image progressive pattern that can be inferred from the incomplete matrix (problem statement), it is considered a correct answer selection[41]. Therefore, a successful RPM discriminator is expected to be able to identify the rationality of the progression patterns within a complete $3\times 3$ progressive matrix.
To address this challenge and enhance the ability to discern underlying patterns, we have devised a novel method. We propose Lico-Net (link-conception net), a network specifically designed for RPM (Raven’s Progressive Matrices) reasoning engines. The core function of Lico-Net is to evaluate the consistency of progressive patterns within a 3×3 RPM matrix and quantify this consistency with a score. By sequentially applying each option provided in an RPM case to complete the matrix, we can utilize Lico-Net to assess the resulting progressive consistency and assign scores to each option accordingly. This approach enables us to efficiently reason through and solve RPM problems. The following is an overview of the detailed design of Lico-Net.
The macro architecture of the Lico-Net we designed is presented in Figure 4. As can be seen from the figure, Lico-Net is divided into four core modules: the Perception Module, the Inference Module, the Concept Fuzzification Module, and the Concept Consistency Evaluation Module. Figure 4 intuitively illustrates our concept; we plan to use the Perception Module to convert all images in the RPM matrix into corresponding representations. Subsequently, the Inference Module and the Concept Fuzzification Module will explore the concepts implicit in the representation matrices with missing entries and depict these concepts using fuzzy representations. Finally, through the Concept Consistency Measurement Module, we will accurately assess the consistency of these concepts. The specific design of the four modules is as follows:
@@ V-A1 Perceptron module of Lico-Net
Aligned with the consensus among RPM solvers on the importance of multi-viewpoint feature extraction [31, 26, 27, 37], Lico-Net employs the Vision Transformer (ViT) for feature extraction, retaining the entire output vectors, which are attention results, as multi-viewpoint features. We denote the number of viewpoints as $L$. Lico-Net then equally processes each viewpoint of the extracted feature. The perceptron of Lico-Net is illustrated in Figure 5.
@@
V-A2 Inference module and Concept Fuzzification Module of Lico-Net
The inference module of Lico-Net processes the representation matrices generated by the perception module for each viewpoint equally and in parallel, resulting in $L$ sets of concepts represented in the form of distributions. During the concept extraction process, the inference module enumerates all possible ways in which the representation matrix can have missing entries, then extracts the concepts implied by each incomplete matrix and represents these concepts as distributions. Ultimately, a distribution distance measurement function is utilized to calculate the distances between concepts within each set, serving as a measure of conceptual consistency. We have illustrated in Figure 6 how the inference module handles a specific missing entry configuration in the representation matrix and transforms it into a concept represented by a distribution.
Figure 6: The process by which the inference module handles the third incomplete representation matrix to extract concepts.
On the left side of the figure 6, it is easy to see that in the Inference module, the MLP with a bottleneck structure extracts the remaining row and column information from the incomplete matrix, which is then fed into a standard Transformer-Encoder to extract the concepts present in the incomplete matrix. The figure shows that we use the Transformer-Encoder to extract 5 feature vectors from 4 row and column informations to represent the concepts, a kind of asymmetric token transformation that relies on learnable vector embeddings, which has been used in ViT. On the right side of the figure, the Concept Fuzzification Module is depicted, illustrating the method of converting vector-form concepts into fuzzy concepts. This involves using a fuzzification head to double the dimensionality of the vectors, then splitting them and treating each part as the mean ($\mu$) and the logarithm of the variance ($\log\sigma^{2}$) of a Gaussian distribution. Subsequently, the reparameterization trick is employed to obtain fuzzy representations of these concepts based on the encoded mean ($\mu$) and variance ($\sigma^{2}$). In this design, both the concept’s baseline ($\mu$) and the degree of fuzziness ($\sigma^{2}$) are learned autonomously. The design of the fuzzification head is shown in Figure 7.
Figure 7: The figure shows that the structure of fuzzification head.

### v16

@@ Abstract
This paper is dedicated to addressing the challenge of enhancing the abstract reasoning capabilities of AI, particularly for tasks involving complex human concepts. We introduce Lico-Net, a novel reasoning engine grounded in deep learning theory, which encodes the logical structure of Raven’s Progressive Matrices (RPM) problems into probabilistic representations. Lico-Net excels in solving RPM tasks. Furthermore, we propose Lico-Net-Bongard, a tailored version of Lico-Net for the Bongard-Logo problem, which also achieves high reasoning accuracy through probabilistic representations. However, we observe a mismatch between the way deep learning algorithms and humans induce reasoning concepts, primarily attributed to the inadequate quality of negative samples. Improper configuration of negative samples can convey erroneous conceptual information to deep learning algorithms, thereby distorting their learning objectives. To address this issue, we propose two novel approaches: first, treating different sample points within reasoning problems as mutual negative samples to alter the existing negative sample structure in the data; second, designing a negative sample generator based on a step-wise linear attention mechanism to produce high-quality negative samples. Experimental results demonstrate that these methods significantly improve the performance of Lico-Net (-Bongard) and other baseline models on the RPM and Bongard-Logo datasets, as well as in the domain of foundational vision model processing, particularly when addressing the NICO dataset’s distribution shift problem. Our findings emphasize the importance of improving negative sample quality for enhancing the abstract reasoning capabilities of deep learning algorithms and suggest that systems represent a promising direction for future research in this field.
@@
Abstract Reasoning, Raven’s Progressive Matrices, Bongard-logo, NICO, Non-Independent and Identically Distributed, Adversarial Learning.
@@ I Introduction
Deep learning, mimicking the human brain, has excelled in numerous fields, particularly in generative tasks like image, text, and video creation through technologies[1] such as Generative Adversarial Networks (GANs) [6], Variational Autoencoders (VAEs) [7], autoregressive models, and Transformer [4]. It has also achieved significant success in computer vision [2, 3], natural language processing [4, 5], generative model construction [6, 7, 8], visual question answering systems [9, 10], and visual abstract reasoning [15, 11, 12]. However, challenges remain in discriminative tasks, where deep learning struggles to fully understand human concepts embedded in the data despite some achievements[13].
This dilemma arises from deep learning models’ over-reliance on superficial statistical patterns, neglecting the deeper structures and logical relationships within human concepts. Despite the tremendous success of deep learning, a significant gap persists in understanding and mimicking complex, ambiguous human concepts, especially in data-rich scenarios[14]. Our research will focus on discriminative issues involving deep and complex human concepts, aiming to explore the root causes of models’ poor performance and develop effective solution, thereby facilitating new advancements in deep learning’s understanding and simulation of human concepts.
@@ II Problem description
Traditional visual discriminative tasks assess the generalization and feature recognition capabilities of deep learning algorithms by evaluating their ability to summarize pixel arrangements in image data. However, in this section, we introduce several non-traditional advanced visual discriminative tasks, which effectively measure the algorithms’ ability to learn complex and abstract patterns underlying visible features.
@@ II-A Raven’s Progressive Matrices
The Raven’s Progressive Matrices (RPM) problem[15] serves as a prestigious benchmark for assessing advanced discriminative capabilities of deep learning algorithms, as it necessitates abstract reasoning, pattern recognition, and problem-solving skills. As a high-level non-verbal task, RPM requires algorithms to possess the ability to generalize and transfer knowledge, thereby challenging them to exhibit cognitive functions akin to human intelligence. Its standardized nature and high correlation with human intelligence measurements render RPM a relevant and consistent test for evaluating the complexity of machine learning models, underscoring its authority in gauging the intelligence level of deep learning algorithms.
@@ II-A1 RAVEN and PGM
The RAVEN database is a collection of Ravens Progressive Matrices problems, each consisting of 16 images. Eight of these images, known as the “statement,” are the main challenge, while the other eight form a pool for option. The goal is to choose the right images to complete a $3\times 3$ matrix with a progressive geometric pattern, following an abstract “rule.” The left side of Figure 1 shows an example of RAVEN. Similarly, PGM provides eight images for the problem statement and options, but its “rule” applies to both rows and columns. The right side of Figure 1 gives a typical example of PGM.
@@ II-B Bongard problem
The Bongard problem[11, 12] is a challenging abstract visual reasoning task that evaluates pattern recognition using minimal training examples, presenting images that follow one of two contrasting rules. It requires abstraction and learning from limited data, posing a significant challenge for both humans and machine learning algorithms. Despite its recognition in AI and cognitive science, there is limited literature on automatically solving these problems, which involve higher cognitive functions like abstract reasoning and concept learning, where machines significantly lag behind humans.
@@ II-B1 Bongrad-logo
The Bongard-Logo[12] problem extends the Bongard problem, serving as a benchmark for human-level concept learning and reasoning. It captures three key aspects of human cognition: context-dependent, analogical, and sample-efficient perception with vast vocabulary. Using the LOGO language, it generates thousands of instances programmatically, challenging standard image classification methods. Experimental results show that deep learning methods fall short, highlighting the need for a more human-like perception. Bongard-Logo problems guide research towards a general visual reasoning architecture that aims to surpass this benchmark.
@@
Bongard-logo problems can be categorized into three parts based on concept domains: 1) Free-form problems (FF), featuring shapes made of random action strokes; 2) Basic shape problems (BA), recognizing pattern of shape combinations; 3) High-level concept problems (HD), assessing abstract concept reasoning, like convexity and symmetry; 4) Combinatorial Abstract Shape (CM) evaluates logical reasoning by testing systems’ ability to recognize new attribute combinations not encountered in training; 5) The Novel Abstract Shape (NV) uses attribute withholding to assess innovative potential, testing models’ capacity to infer new rules when an attribute and its combinations are excluded from training.
@@ III-A RPM solver
Image reasoning models such as CoPINet[21], LEN+teacher[20], and DCNet[22] focus on learning disparities and rules. Meanwhile, NCD[24], SCL[25], SAVIR-T[26], and neural symbolism systems (PrAE, NVSA, ALANS[27, 28, 29]) enhance both interpretability and accuracy. RS-CNN and RS-TRAN[30] excel in solving RPM problems, while Triple-CFN[37] stands out by implicitly extracting and indexing concept and reasoning unit info, boosting reasoning accuracy. CRAB[31], based on Bayesian modeling, has made strides with a tailored conservatory for the RAVEN database, though core challenges remain untackled, leaving the scientific community eager for the broader impact of these innovative methods.
@@ III-B Bongard solver
Addressing Bongard problems, approaches include language feature models, CNNs, and synthetic data. Depweg et al.[11] used Bayesian reasoning with formal language but faced scalability issues. Kharagorgiev and Yun[35] focused on feature extraction, and Nie et al.[12] experimented with CNNs and meta-learning with limited success. PMoC[36] offers a probabilistic approach to assess group sample likelihoods, proving effective. Triple-CFN[37] stands out for its consistent effectiveness on both Bongard-Logo and RPM problems.
@@ III-C NICO solver
Stable learning [33] addresses O.o.D. issues by reweighting samples to correct domain shifts and reduce feature correlations, boosting model performance on unseen tests. DecAug [34] tackles the 2D O.o.D. problem in NICO data by adding domain labels and splitting the model into domain and object learning branches with orthogonal parameter updates for information decoupling, albeit at the cost of increased workload and categorization difficulty. SimCLR [43] enhances representation learning with data augmentation, a learnable nonlinear projector, and the InfoNCE loss. VICReg [42] introduces three regularization terms on embedding variance for self-supervised image representation learning.
@@ IV Methodology
In this study, we highlight the importance of deep learning algorithms learning probabilistic representations of complex human concepts. While probabilistic representations enhance performance, we observe a discrepancy between how algorithms and humans summarize concepts and logical patterns. The core issue lies in the inadequate configuration of negative samples, which conveys erroneous information and distorts learning objectives. Therefore, this study posits that, in addition to striving to develop deep learning algorithms capable of effectively learning from low-quality negative samples, improving the quality of negative samples is also a crucial strategy.
@@ V Discriminator of Concepts (DC)
This section introduces Lico-Net, a novel reasoning engine built upon Deep Neural Networks (DNNs), designed to tackle the Raven’s Progressive Matrices (RPM) problem. By distilling complex human reasoning concepts into probabilistic representations, Lico-Net demonstrates exceptional reasoning performance. We further propose Lico-Net-Bongard, a model specifically designed for solving the Bongard Problems, which achieves outstanding reasoning accuracy through explicit probabilistic representation learning. Preliminary experiments on these models have uncovered a concept mismatch phenomenon, highlighting the limitations of deep models in learning implicit concepts from data. Additionally, using the NICO dataset, we demonstrate how variations in context distribution can undermine the models’ ability to learn these concepts.
@@ V-A A novel baseline, Lico-Net, for RPM problems
The RPM problem requires participants to select an answer from an option pool to complete a $3\times 3$ progressive matrix of images. If the completed matrix follows a specific progressive pattern inferable from the problem statement, the answer is considered correct [41]. Therefore, a successful RPM discriminator should be able to recognize the consistency of various progressive patterns within a complete $3\times 3$ matrix. To address this challenge and enhance the ability to discern underlying patterns, we have devised a novel approach. We propose Lico-Net (Link-Conception Network), whose core function is to assess the consistency of progressive patterns within a $3\times 3$ RPM matrix and quantify this consistency with a score. By sequentially applying each option provided in an RPM case to the matrix, we can utilize Lico-Net to evaluate the resulting progressive consistency and assign scores to each option accordingly. This method enables us to efficiently reason through and solve RPM problems. The following is an overview of the detailed design of Lico-Net.
The macro architecture of the Lico-Net we designed is presented in Figure 4. As can be seen from the figure, Lico-Net is divided into four core modules: the Perception Module, the Reasoning Module, the Concept Extraction Module, and the Concept Consistency Evaluation Module. Figure 4 intuitively illustrates our concept; we plan to use the Perception Module to convert all images in the RPM matrix into corresponding representations. Subsequently, the Reasoning Module and the Concept Extraction Module will explore the concepts implicit in the representation matrices with missing entries and depict these concepts using probabilistic representations. Finally, through the Concept Consistency Measurement Module, we will accurately assess the consistency of these concepts. The specific design of the four modules is as follows:
@@ V-A1 Perceptron Module of Lico-Net
Aligned with the consensus among RPM solvers on the importance of multi-viewpoint feature extraction [30, 25, 26, 37, 32], Lico-Net employs the Vision Transformer (ViT) for feature extraction, retaining the entire output vectors, which are attention results, as multi-viewpoint features. We denote the number of viewpoints as $L$. Lico-Net then equally processes each viewpoint of the extracted feature. The perceptron of Lico-Net is illustrated in Figure 5.
@@
V-A2 Reasoning Module and Concept Extraction Module of Lico-Net
The Reasoning Module and the Concept Extraction Module of Lico-Net process the representation matrices generated by the Perception Module for each viewpoint equally and in parallel, resulting in $L$ sets of concepts presented in probabilistic representations. Specifically, the Reasoning Module enumerates all possible ways in which the representation matrix can have missing entries, then extracts the concepts implied by each incomplete matrix. The Concept Extraction Module subsequently represents these concepts as probabilistic representations. We have illustrated this process in Figure 6.
Figure 6: The process by which the Reasoning Module handles the third incomplete representation matrix to extract concepts.
  1. 1.
    On the left side of the figure 6, it is easy to see that in the Reasoning Module, the MLP with a bottleneck structure extracts the remaining row and column information from the incomplete matrix, which is then fed into a standard Transformer-Encoder to extract the concepts present in the incomplete matrix. The figure shows that we use the Transformer-Encoder to extract five feature vectors from four row and column information to represent the concepts. The Transformer-Encoder capable of achieving asymmetric transformation in token counts is realized by incorporating learnable class vectors, and this technique has been successfully applied in the Vision Transformer (ViT) [19].
  2. 2.
    On the right side of the figure, the Concept Extraction Module is depicted, illustrating the method of converting vector-form concepts into probabilistic representation. This involves using a ‘probabilistic head’ to double the dimensionality of the vector-form concepts, then splitting them and treating each part as the mean ($\mu$) and the logarithm of the variance ($\log\sigma^{2}$) of a Gaussian distribution. This achieves the transformation of vector-form concepts into distribution-form concepts. Subsequently, the reparameterization trick[7] is employed to obtain probabilistic representations of these distribution-form concepts. In this design, both the concept’s baseline ($\mu$) and the degree of uncertainty ($\sigma^{2}$) are learned autonomously. The design of the probabilistic head is shown in Figure 7.
Figure 7: The figure shows that the structure of probabilistic head.